X=reaction time, Y=length of life of electronic components Continuous joint probability distribution:

Compute: P(0 < X < 0.6 and 0.25 < Y < 0.5)



1
Expert's answer
2021-11-16T16:20:18-0500

"f(x,y)=\left\{\begin{matrix}

4xy, 0<x<1.0, 0<y<1 & \\

0, elsewhere &

\end{matrix}\right."

"P= \int \int F(x,y) dx dy"

P(0<x<0.5 + 0.25<y<0.5) fall in this region.

"P = \int^{0.6}_0 \int^{0.5}_{0.25} 4 xy dydx \\

P = \int^{0.6}_0 [4x (\frac{y^2}{2})]^{0.5}_{0.25} dx \\

= \int^{0.6}_0 2x [(\frac{1}{2})^2 -(\frac{1}{4})^2]dx \\

= (\frac{3}{8})\int^{0.6}_0 xdx \\

= \frac{3}{8}(\frac{x^2}{2})^{0.6}_0 \\

= \frac{3}{8}(0.6)^2 \\

= 0.0675"